# Organizing Unstructured Image Collections using Natural Language

### v1

@@ B.1 Benchmark Construction of COCO-4c and Food-4c
Dataset Basic Criterion Number of Classes
COCO-4c Activity 64
@@
Food-4c Food Type 101
@@
B.2 Details on Hard Grouping Criteria Annotation
Table 6: Annotated criteria for the six benchmarks. The basic criteria are annotated on per-image level for each benchmark, while the hard criteria (those not in the basic criteria) are further exhaustively annotated by human annotators for further evaluating the performance of the rule proposer in SMC task.
@@ B.2 Details on Hard Grouping Criteria Annotation
While we have established more rigorous and challenging benchmarks such as COCO-4c and Food-4c, which feature up to four distinct grouping criteria, these annotated criteria sets do not encompass all potential grouping criteria within the image collections. This is particularly true for more complex and realistic datasets like COCO-4c, Food-4c, and Action-3c. As a result, the performance differences between different criteria proposers on these basic criteria, as shown in Fig. 4, tend to be close to each other, limiting our understanding of each proposer’s ability to generate comprehensive grouping criteria.
@@
(2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g.,, Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
@@
Appendix C Further Implementation Details
In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App. C.1 and the Semantic Grouper in App. C.2.
C.1 Prompts and Implementation Details of Criteria Proposer
@@
In Tab. 7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2024). Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images. Each subset is then stitched into an $8\times 8$ image grid, treated as a single image, and fed into the MLLM. For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab. 7. After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set. Finally, we deduplicate the discovered criteria and accumulate them into the criteria pool.
@@
In Tab. 8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024b). For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image. These tags are selected from the WordNet (Miller, 1995) vocabulary, which contains 118k English synsets, and represent the semantic content of the images. We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging. Next, we embed the assigned tags into the prompt shown in Tab. 8 to carry the semantics of the entire image set and query the LLM to propose grouping criteria. The criteria proposed by the LLM are then added to the criteria pool. Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g.,, 128k for Llama-3.1-8B) for the datasets used in our experiments. However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
@@
We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab. 9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024b) in Tab. 10. Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language. The generated captions are then randomly shuffled and split into disjoint subsets, each containing 400 captions. Next, we embed the captions from each subset into the prompt shown in Tab. 10 and use it to query the LLM to propose grouping criteria for the images represented by the captions. After iterating through all subsets, we take the union of the proposed criteria across subsets as the discovered criteria for the target image set. Finally, we deduplicate these criteria and add them to the criteria pool. Due to the context window limitations of LLMs, embedding all captions into a single prompt is infeasible. To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset. This strategy allows us to remain within the context length limits of modern LLMs (e.g.,, 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
@@
In Tab. 11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024b). Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab. 11 and ask the LLM to merge similar criteria and rephrase their names to enhance clarity. This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
C.2 Prompts and Implementation Details of Semantic Grouper
@@
Appendix D Further Implementation Details of the Compared Methods
@@
Implementation details of MMaP (Yao et al., 2024): We closely followed the training configuration outlined in the original MMaP paper. Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset. We then prompt-tuned CLIP-ViT/B32 using Adam with a momentum of 0.9, training the model for 1,000 epochs for each criterion across all datasets. Hyperparameters were optimized according to the loss score of MMaP, with the learning rate searched in {0.1, 0.05, 0.01, 0.005, 0.001, 0.0005}, weight decay in {0.0005, 0.0001, 0.00005, 0.00001, 0}, $\alpha$ and $\beta$ in {0.0, 0.1, 0.2, …, 1.0}, and $\lambda$ fixed at 1 for all experiments. After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
Appendix E Supplementary Results of the Quantitative Experiments
In this section, we present additional numerical experiment results to supplement the figures in the main paper. In Sec. E.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework. In Sec. E.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks. Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods. Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in Sec. E.4.
E.1 Supplementary Results for Criteria Proposer Evaluation
We provide detailed numerical results corresponding to Fig. 4(a) in Tab. 20 and Fig. 4(c) in Tab. 21 for the six tested benchmarks.
Although captions generated by the MLLM may exhibit some information loss (e.g.,, ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g.,, introducing objects not present in the images) Liu et al. (2024a), these issues generally occur at the object or fine-grained attribute level. However, when reasoning about grouping criteria for SMC task, the focus is on identifying general thematic elements shared across the image set. As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
Table 20: Comparison of True Positive Rate (TPR) (%) for criteria proposers across the six SMC benchmarks. TPR performance is reported for both Basic and Hard ground-truth criteria. The best performance is highlighted in bold.
@@ E.1 Supplementary Results for Criteria Proposer Evaluation
Table 21: Study of the impact of data scale on criteria discovery. The Caption-based Proposer is used for criteria discovery, and TPR performance (%) is reported on the Hard ground-truth criteria sets across the six SMC benchmarks for different data scales. The best performance is highlighted in bold.
@@
E.2 Supplementary Results for Semantic Grouper Evaluation
In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Fig. 5. Specifically, we provide detailed results for the evaluation of the six tested datasets as follows:
  • COCO-4c (Fig. 5(a)) in Tab. 22
  • Card-2c (Fig. 5(b)) in Tab. 23
  • Action-3c (Fig. 5(c)) in Tab. 24
  • Food-4c (Fig. 5(d)) in Tab. 25
  • Fruit-2c (Fig. 5(e)) in Tab. 26
  • Clevr-4c (Fig. 5(f)) in Tab. 27
Table 22: Comparison of Semantic Groupers on COCO-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(a).
@@ E.2 Supplementary Results for Semantic Grouper Evaluation
Table 23: Comparison of Semantic Groupers on Card-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(b).
@@
Table 24: Comparison of Semantic Groupers on Action-3c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(c).
@@
Table 25: Comparison of Semantic Groupers on Food-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(d).
@@
Table 26: Comparison of Semantic Groupers on Fruit-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(e).
@@
Table 27: Comparison of Semantic Groupers on Clevr-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(f).
@@
E.3 Supplementary Results for Comparison Criteria-Conditioned Clustering Methods
@@
CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
COCO-4c Activity 62.6 73.5 51.3 53.2 33.8 - 44.1 48.9
Location 34.3 51.5 58.5 54.0 35.3 - 55.2 55.6
Mood 22.4 43.3 23.2 40.4 20.9 - 38.1 32.6
Time of Day 40.6 74.1 62.8 65.2 45.7 - 67.6 56.7
$Avg.$ 40.1 60.6 48.9 53.2 33.9 - 51.2 48.4
Food-4c Food Type 90.6 94.6 36.0 52.6 48.9 - 34.6 54.2
Cuisine 54.9 81.4 46.8 42.4 31.7 - 47.0 65.9
Course 63.5 84.7 70.5 89.5 48.6 - 69.1 85.7
Diet 47.6 59.9 48.5 62.1 45.9 - 41.5 54.0
$Avg.$ 64.1 80.2 50.5 61.7 43.8 - 48.1 64.9
Clevr-4c Color 77.7 94.0 51.2 43.2 75.3 - 70.3 63.4
Texture 34.1 41.9 64.9 26.4 56.5 - 65.3 42.1
Count 43.7 81.5 46.9 39.0 53.9 - 65.7 73.3
Shape 71.1 72.7 70.0 38.7 65.5 - 58.4 38.5
$Avg.$ 56.7 72.5 58.3 36.8 62.8 - 64.9 54.3
Action-3c Action 97.1 99.2 86.4 58.7 51.3 - 82.8 76.3
Location 66.7 67.1 82.0 52.9 59.4 - 69.8 55.2
Mood 75.5 80.7 60.8 57.4 71.0 - 52.3 50.2
$Avg.$ 79.8 82.3 76.4 56.3 60.6 - 68.3 60.6
Card-2c Suit 47.9 69.5 54.9 65.6 41.3 - 54.5 73.6
Rank 35.0 64.2 94.6 96.8 32.6 - 92.1 95.1
$Avg.$ 41.4 66.9 74.8 81.2 36.9 - 73.3 84.3
Fruit-2c Species 84.0 93.1 69.3 66.9 58.8 - 76.9 70.7
Color 54.8 83.5 57.2 43.3 43.3 - 53.3 51.5
$Avg.$ 69.4 88.3 63.3 55.1 51.0 - 65.1 61.1
E.4 Supplementary Results for Studying the Necessity of Multi-granularity Cluster Generation.
We present expanded results in Tab. 29 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Fig. 6 in the main paper.
Table 29: Ablation study of multi-granularity refinement on the six SMC benchmarks. We compare three ways of constructing cluster names: Initial Names (IN), Flat Refinement (FR), Multi-granularity Refinement (MR). We report Clustering Accuracy (CAcc) and Semantic Accuracy (SAcc)as percentages (%). Average (Avg.) CAcc and SAcc across different criteria on each dataset is also provided. These expanded results correspond to the plotting shown in Fig. 6.

### v4

@@ C.1 Construction of COCO-4c and Food-4c
Dataset Number of Images Basic Criterion Number of Classes
COCO-4c 5,000 Activity 64
@@
Food-4c 25,250 Food Type 101
@@
C.2 Details on Hard Grouping Criteria Annotation
Table 6: Annotated criteria for the six benchmarks. The basic criteria are annotated on per-image level for each benchmark, while the hard criteria (those not in the basic criteria) are further exhaustively annotated by human annotators for further evaluating the performance of the rule proposer in OpenSMC task.
@@ C.2 Details on Hard Grouping Criteria Annotation
While we have established more rigorous and challenging benchmarks such as COCO-4c and Food-4c, which feature up to four distinct grouping criteria, these annotated criteria sets do not encompass all potential grouping criteria within the image collections. This is particularly true for more complex and realistic datasets like COCO-4c, Food-4c, and Action-3c. As a result, the performance differences between different criteria proposers on these basic criteria, as shown in Figure 4, tend to be close to each other, limiting our understanding of each proposer’s ability to generate comprehensive grouping criteria.
@@
(2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g., Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
@@
Appendix D Further Details of Evaluation Protocol
Further Discussion on Clustering Accuracy (CAcc). Clustering Accuracy (CAcc) (Han et al., 2021) is evaluated by applying the Hungarian algorithm (Kuhn, 1955) to determine the optimal assignment between the predicted cluster indices and ground-truth labels. As extensively discussed in the GCD (Vaze et al., 2022) literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation. On the other hand, if the number of predicted clusters is lower than the number of ground-truth classes, the extra classes are assigned to a null set, and all instances with those ground-truth labels are similarly considered incorrect. Thus, CAcc is maximized only when the number of predicted clusters matches the number of ground-truth clusters.
[…]
Appendix E Further Implementation and Prompt Details
In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App. E.1 and the Semantic Grouper in App. E.2.
E.1 Further Details of Criteria Proposer
@@
In Tab. 7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2025). Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images. Each subset is then stitched into an $8\times 8$ image grid, treated as a single image, and fed into the MLLM. For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab. 7. After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set. Finally, we deduplicate the discovered criteria and accumulate them into the criteria pool.
@@
In Tab. 8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024a). For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image. These tags are selected from the WordNet (Miller, 1995) vocabulary, which contains 118k English synsets, and represent the semantic content of the images. We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging. Next, we embed the assigned tags into the prompt shown in Tab. 8 to carry the semantics of the entire image set and query the LLM to propose grouping criteria. The criteria proposed by the LLM are then added to the criteria pool. Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g., 128k for Llama-3.1-8B) for the datasets used in our experiments. However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
@@
We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab. 9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024a) in Tab. 10. Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language. The generated captions are then randomly shuffled and split into disjoint subsets, each containing 400 captions. Next, we embed the captions from each subset into the prompt shown in Tab. 10 and use it to query the LLM to propose grouping criteria for the images represented by the captions. After iterating through all subsets, we take the union of the proposed criteria across subsets as the discovered criteria for the target image set. Finally, we deduplicate these criteria and add them to the criteria pool. Due to the context window limitations of LLMs, embedding all captions into a single prompt is infeasible. To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset. This strategy allows us to remain within the context length limits of modern LLMs (e.g., 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
@@
Since the Proposers operate on subsets, the criteria accumulated in $\mathcal{R}$ may include redundant or noisy entries. Therefore, we further include a criteria poop refinement process. Specifically, we input all initially proposed criteria from $\mathcal{R}$ into a LLM, prompting it to consolidate similar criteria and discard noisy ones. This process refines and updates the criteria pool $\mathcal{R}$ for next stage.
In Tab. 11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024a). Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab. 11 and ask the LLM to merge similar criteria and rephrase their names to enhance clarity. This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
E.2 Further Details of Semantic Grouper
@@
Appendix F Further Details of the Compared Methods
@@
[…]
Implementation details of MMaP and MSub (Yao et al., 2024, 2025): We closely followed the training configuration outlined in the original MMaP and MSub paper. Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset. We then prompt-tuned CLIP-ViT/B32 using Adam with a momentum of 0.9, training the model for 1,000 epochs for each criterion across all datasets. Hyperparameters were optimized according to the loss score of MMaP, with the learning rate searched in {0.1, 0.05, 0.01, 0.005, 0.001, 0.0005}, weight decay in {0.0005, 0.0001, 0.00005, 0.00001, 0}, $\alpha$ and $\beta$ in {0.0, 0.1, 0.2, …, 1.0}, and $\lambda$ fixed at 1 for all experiments. After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
Appendix G Further Quantitative Experimental Results
In this section, we present additional numerical experiment results to supplement the figures in the main paper. In § G.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework. In § G.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks. Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods. Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § G.4.
G.1 Further Results for Criteria Proposer Study
We provide detailed numerical results corresponding to Figure 4 in Tab. 20 and Figure 5 in Tab. 21 for the six tested benchmarks.
Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g., introducing objects not present in the images) Liu et al. (2024a), these issues generally occur at the object or fine-grained attribute level. However, when reasoning about grouping criteria for OpenSMC task, the focus is on identifying general thematic elements shared across the image set. As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
Table 20: Comparison of True Positive Rate (TPR) (%) for criteria proposers across the six OpenSMC benchmarks. TPR performance is reported for both Basic and Hard ground-truth criteria. The best performance is highlighted in bold.
@@ G.1 Further Results for Criteria Proposer Study
Table 21: Study of the impact of data scale on criteria discovery. The Caption-based Proposer is used for criteria discovery, and TPR performance (%) is reported on the Hard ground-truth criteria sets across the six OpenSMC benchmarks for different data scales. The best performance is highlighted in bold.
@@
G.2 Further Results for Semantic Grouper Study
In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Figure 6. Specifically, we provide detailed results for the evaluation of the six tested datasets as follows:
In addition, we present the statistics of the predicted clusters at each granularity level in Tab. 28.
Table 22: Comparison of Semantic Groupers on COCO-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(a).
@@ G.2 Further Results for Semantic Grouper Study
Table 23: Comparison of Semantic Groupers on Card-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(b).
@@
Table 24: Comparison of Semantic Groupers on Action-3c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(c).
@@
Table 25: Comparison of Semantic Groupers on Food-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(d).
@@
Table 26: Comparison of Semantic Groupers on Fruit-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(e).
@@
Table 27: Comparison of Semantic Groupers on Clevr-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Figure 6(f).
@@
Table 28: Summary of cluster counts across six benchmarks for the comparison of semantic groupers. The results yield by the main Caption-based Grouper is reported. Specifically, we report: i) GT: the number of ground-truth clusters; ii) Pred-Init: predicted clusters from initial names; iii) Pred-Coarse: predicted coarse-grained clusters after multi-granularity refinement; iv) Pred-Middle: predicted middle-grained clusters after multi-granularity refinement; and v) Pred-Fine: predicted fine-grained clusters after multi-granularity refinement.
Dataset Criteria GT Pred-Init Pred-Corase Pred-Middle Pred-Fine
COCO-4c Activity 64 203 12 23 52
Location 19 145 7 14 28
Mood 20 122 15 25 30
Time of Day 6 96 2 8 31
Food-4c Food Type 101 301 7 37 127
Cuisine 15 141 9 18 53
Course 5 97 4 12 78
Diet 4 139 5 8 64
Action-3c Action 40 71 8 15 51
Location 10 82 5 10 67
Mood 4 95 6 18 55
Clevr-4c Color 10 25 6 12 17
Texture 10 23 2 5 12
Shape 10 22 5 11 14
Count 10 11 2 4 11
Card-2c Rank 14 147 4 7 16
Suit 5 56 4 7 30
Fruit-2c Species 34 54 8 25 38
Color 15 66 5 15 39
G.3 Further Comparative Results with TCMC Methods
@@
CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
COCO-4c Activity 62.6 73.5 51.3 53.2 44.0 52.1 33.8 - 35.9 - 44.1 48.9
Location 34.3 51.5 58.5 54.0 51.2 52.9 35.3 - 37.4 - 55.2 55.6
Mood 22.4 43.3 23.2 40.4 15.9 39.3 20.9 - 23.0 - 38.1 32.6
Time of Day 40.6 74.1 62.8 65.2 55.5 64.1 45.7 - 47.8 - 67.6 56.7
$Avg.$ 40.1 60.6 48.9 53.2 41.6 52.1 33.9 - 36.0 - 51.2 48.4
Food-4c Food Type 90.6 94.6 36.0 52.6 33.1 46.5 48.9 - 52.4 - 34.6 54.2
Cuisine 54.9 81.4 46.8 42.4 43.9 36.3 31.7 - 35.2 - 47.0 65.9
Course 63.5 84.7 70.5 89.5 67.6 83.4 48.6 - 52.1 - 69.1 85.7
Diet 47.6 59.9 48.5 62.1 45.6 56.0 45.9 - 49.4 - 41.5 54.0
$Avg.$ 64.1 80.2 50.5 61.7 47.5 55.5 43.8 - 47.3 - 48.1 64.9
Clevr-4c Color 77.7 94.0 51.2 43.2 47.8 44.0 75.3 - 84.7 - 70.3 63.4
Texture 34.1 41.9 64.9 26.4 61.5 27.2 56.5 - 65.9 - 65.3 42.1
Count 43.7 81.5 46.9 39.0 43.5 39.8 53.9 - 63.3 - 65.7 73.3
Shape 71.1 72.7 70.0 38.7 66.6 39.5 65.5 - 74.9 - 58.4 38.5
$Avg.$ 56.7 72.5 58.3 36.8 54.8 37.6 62.8 - 72.2 - 64.9 54.3
Action-3c Action 97.1 99.2 86.4 58.7 88.1 55.3 51.3 - 55.0 - 82.8 76.3
Location 66.7 67.1 82.0 52.9 83.7 49.5 59.4 - 63.1 - 69.8 55.2
Mood 75.5 80.7 60.8 57.4 62.5 54.0 71.0 - 74.7 - 52.3 50.2
$Avg.$ 79.8 82.3 76.4 56.3 78.1 52.9 60.6 - 64.3 - 68.3 60.6
Card-2c Suit 47.9 69.5 54.9 65.6 47.5 60.7 41.3 - 44.0 - 54.5 73.6
Rank 35.0 64.2 94.6 96.8 87.2 91.9 32.6 - 35.3 - 92.1 95.1
$Avg.$ 41.4 66.9 74.8 81.2 67.3 76.3 36.9 - 39.6 - 73.3 84.3
Fruit-2c Species 84.0 93.1 69.3 66.9 68.1 58.6 58.8 - 62.2 - 76.9 70.7
Color 54.8 83.5 57.2 43.3 56.0 35.0 43.3 - 46.7 - 53.3 51.5
$Avg.$ 69.4 88.3 63.3 55.1 62.0 46.8 51.0 - 54.4 - 65.1 61.1
G.4 Further Results for Multi-granularity Clustering Study.
We present expanded results in Tab. 30 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Figure 7 in the main paper.
Table 30: Ablation study of multi-granularity refinement on the six OpenSMC benchmarks. We compare three ways of constructing cluster names: Initial Names (IN), Flat Refinement (FR), Multi-granularity Refinement (MR). We report Clustering Accuracy (CAcc) and Semantic Accuracy (SAcc)as percentages (%). Average (Avg.) CAcc and SAcc across different criteria on each dataset is also provided. These expanded results correspond to the plotting shown in Figure 7.